OpenStudy (anonymous):

7 - (2x + 11) + 3(3 - x) = 20

OpenStudy (anonymous):

\(\Large 7 - (2x + 11) + 3(3 - x) = 20\) \(\Large 7-2x-11+9-3x=20\) \(\Large -5x+5=20\) \(\Large 5x=5-20=-15\) \(\Large x=\frac{-15}{5}=-3\)